How old is Bobby if, in 4 years, he will be 3 times as old as he is today?"
I got 3(B+4)THUS B=12
I don't think I got this correct though

Found 2 solutions by Alan3354, macston:
Answer by Alan3354(69443)

How old is Bobby if, in 4 years, he will be 3 times as old as he is today?"
I got 3(B+4)THUS B=12
Check and see.
Does 12+4 = 3*12 ?
--------------
B + 4 = 3B
2B = 4
B = 2

Answer by macston(5194)

Let B=Bobby's age
How old is Bobby if, in 4 years (B+4yrs) , he will be 3 times as old as he is today? (3B)"
B+4yrs=3B Subtract B from each side.
4yrs=2B Divide each side by 2
2 yrs=B ANSWER: Bobby is 2 years old
CHECK:
3B=B+4yrs
3(2yrs)=2+4yrs
6yrs=6yrs